# Video-XL: Extra-Long Vision Language Model for Hour-Scale Video Understanding

### v2

@@ 4.3 Ablation studies
Using the Bunny 695k dataset, we trained two models: one without compression and the other with random compression ratios selected from ${2,4,8}$. For the compressed model, we applied varying compression ratios when testing on the video benchmark MLVU and image benchmarks MME and MMBench. The results can be seen in Table 3. Remarkably, even with an 8×8$\times$8 × compression ratio, the compressed models demonstrated promising performance compared to the baseline.
  Model MLVU MME MMBench
  Baseline 57.0 1534 (395) 71.6
2×2$\times$2 × Com. 56.7 1520 (348) 71.4
4×4$\times$4 × Com. 56.5 1518 (328) 71.4
8×8$\times$8 × Com. 56.0 1501 (320) 70.3
@@
Table 3: Ablation studies on the impact of Visual Context Latent Summarization.
4.3.1 Visualization Results
We provide some typical cases of long video understanding tasks, including video summarization, surveillance anomaly detection, and Ad placement identification, as shown in Figure 3.
Figure 3: Visualization results of Video-XL.
@@ 5 Conclusions
We propose Video-XL, an extra-long vision-language model designed for hour-scale video understanding. By adapting the LLM as a visual condenser, we introduce Visual Context Latent Summarization, which condenses visual contexts into highly compact forms. Despite being trained on limited image data, Video-XL achieves promising results on several long video understanding benchmarks and demonstrates its capability in handling long contexts in the Needle-in-a-Haystack evaluation. In the future, we plan to scale up both the training data and model to further enhance performance in long video understanding.

### v3

@@ 4.3 Ablation studies
We conducted ablation studies on the proposed visual compression mechanism and the VICO dataset, as shown in Table 4.
The effectiveness of visual compression. Using the Bunny 695k dataset, we trained two models: one without compression and the other with random compression ratios selected from ${2,4,8,16}$. For the compressed model, we applied varying compression ratios when testing on the video benchmark MLVU and image benchmarks MME and MMBench. Remarkably, even with an 16×16$\times$16 × compression ratio, the compressed models demonstrated promising performance compared to the baseline.
The effectiveness of VICO data. We trained four models using different data: (a) Bunny 695k alone; (b) Bunny 695k combined with NeXTQA 32k; (c) Bunny 695k combined with CinePile 10k; (d) Bunny 695k combined with long video caption 5k (generated similarly to  [18]); and (e) Bunny 695k combined with VICO 5k. Remarkably, even with just 5k VICO data, Video-XL outperforms the one trained on NeXTQA 32k. Additionally, the main event/action ordering task demonstrates greater benefits compared to the captioning task, as it encourages the model to retrieve key clips from long sequences.
     Impact of Visual Compression Impact of VICO
  Model MLVU MME MMBench Train Data Video-MME MLVU
  Baseline 57.0 1534 (395) 71.6 Bunny 52.6 57.0
2×2$\times$2 × Com. 56.7 1520 (348) 71.4 Bunny + NExTQA 32k 53.3 59.9
8×8$\times$8 × Com. 56.4 1515 (326) 71.2 Bunny + CinePile 10k 52.9 59.7
16×16$\times$16 × Com. 56.1 1503 (324) 70.6 Bunny + LongCap 10k 53.2 57.5
@@
Table 4: Ablation studies on the impact of Visual Context Latent Summarization and VICO dataset.
4.3.1 Qualitative Results
We provide some qualitative results of Video-XL in Figure 3, showcasing typical cases of long video understanding. All video samples are longer than 15 minutes. Video-XL demonstrates its ability to identify specific details in long videos, such as anomaly clips in surveillance footage or a placed ad in a video, highlighting its proficiency in retrieving and interpreting visual information over extended video contexts. Additionally, Video-XL can accurately summarize a movie by describing its main characters, surroundings, and events. These capabilities emphasize Video-XL’s potential to process and understand extremely long videos.
Figure 3: Visualization results of Video-XL in some long video understanding tasks.
@@ 5 Conclusions
We propose Video-XL, a VLM designed for hour-scale video understanding. We introduce Visual Context Latent Summarization, which condenses visual contexts into highly compact forms. Video-XL achieves promising results on several long video understanding benchmarks and demonstrates its capability in handling long contexts in the Needle-in-a-Haystack evaluation.